# A Short Proof of the Poincaré Conjecture

### v1

@@ 2. Patterns in a Tetrahedron
A track is a connected pattern. If two patterns $P$ and $Q$ intersect each $1$-simplex in the same number of points then the patterns are said to be equivalent. Two equivalent disjoint tracks in the same $2$-complex are said to be parallel. We investigate tracks and patterns in a tetrahedron $T$, which we regard as the $2$-skeleton $|\rho^{2}|$ of a $3$-simplex $\rho$. We call a track in $T$ an $n$-track if it intersects $n$ edges.
@@
Figure 3 shows a $12$-track. A pattern in $T$ can only have two types of track. There can be $3$-tracks, each of which separates one of the corner vertices from the other three vertices. There can be one other parallel set of tracks each of which is a $4n$-track for the same positive integer $n$. Each such track separates the four vertices into two pairs. If $n>0$ is even, it separates $u,v$ from $w,z$ and if $n>1$ is odd it separates $u,z$ from $v,w$.
@@
Orient a track in $T$ by choosing a positive direction as one goes along the track. At adjacent points of intersection of a track with an edge the directions of the track will be opposite to each other. This gives what we call a + - pair, i.e. a pair of points - not necessarily adjacent - of intersection points of an edge with a track where the track has opposite directions. This will be of more significance for singular “tracks ” or stracks as shown in Figures 4 and 5, as a pair of points of intersection on an edge can be removed by a homotopy if and only if the pair is a + - pair. A track in $T$ has a + - pair if and only if it is not a $3$-track or a $4$-track, since any track which intersects an edge more than once will have a + - pair.
@@
Let $f:S^{2}\rightarrow M$ be a general position map (see Hempel [3], Chapter 1), in which $f$ is in general position with respect to a triangulation $K$ of $M$. An $i$-piece of $f$ is defined to be a component of $f^{-1}(\sigma)$ where $\sigma$ is an $(i+1)$-simplex of $K$. Thus a 00-piece is a point of $S^{2}$. A $1$-piece is either an scc (simple closed curve) or an arc joining two 00-pieces. If there are no $1$-pieces that are scc’s, then each $2$-piece is a disc whose boundary is a union of $1$-pieces. One can use surgery along simple closed curves to change $f$ to a map in which there are no $1$-pieces that are scc’s, and in which every $2$-piece is a disc. The $2$-pieces will then give a cell decomposition (tessellation) of the $2$-sphere.
@@
If $P$ is a $1$-piece with end points $u,v$ whose images under $f$ are in the same $1$-simplex, then the restriction of $f$ to $P$ is called a returning arc.
@@
In [1] there is a mistake on page 253 of that section on simplifying surface maps. It is incorrectly asserted there that any pair of points in the intersection of the boundary of a $2$-piece with a $1$-simplex can be removed by a homotopy. In fact the pair needs to be a + - pair. Suppose $P$ is $2$-piece for which there is a + - pair $u,v$ in the strack $\delta P$ and that $\gamma$ is the $1$-simplex containing this pair. Let $s:S^{1}\rightarrow\delta P$ be as in the definition of a strack. Possibly by choosing a different + - pair in $\gamma$ we can assume that there is an arc $I=[p,q]\subset S^{1}$ such that $s(p)=u$ and $s(q)=v$ and $s(I)$ intersects $\gamma$ only in its end points $u,v$. There will be a map $s^{\prime}:[p,q]\rightarrow P$ which is close to $s$, for which $s^{\prime}(p)=u,s^{\prime}(q)=v$ and the image of the open interval $(p,q)$ is contained in the interior of the $3$-simplex $\rho$ containing $P$. Any two maps of $I$ into the interior of $\rho$ are homotopic. In particular there will be such a map that is close to the interval in $\gamma$ joining $p,q$. By adjusting $f:S^{2}\rightarrow M$ by such a homotopy and then by another homotopy in a neighbourhood of $I$, we can get a new map $f$ in which the pair $u,v$ has been removed. This is illustrated in Figure 4, where it is shown how the removal of a returning arc will create returning arcs in two other simplexes that have the same $1$-face.
@@
Also the pair can be removed without disturbing any other points of intersection with the $1$-skeleton.
@@
Figure 5 shows a 12-strack in $T$ that has no + - pairs of points. I conjecture that a strack in $T$ must intersect all six edges to have a + - pair.
@@
Figure 6 shows a spattern in $T$ that is a union of a red 12-track and a blue 3-track, and its underlying pattern, which is also a $3$-track and a $12$-track.
@@
Given two finite subsets $F_{1},F_{2}$ of the closed interval $[0,1]$ and a bijection $\beta_{I}:F_{1}\rightarrow F_{2}$, there is a continuous map $\phi_{I}:I\rightarrow I$ which restricts to $\beta$ on $F_{1}$ and to the identity on $\{0,1\}$. There is a homotopy between $\phi_{I}$ and the identity map.
@@ 3. The Proof
Let $M$ be a fake 3-ball. Let $M_{0}$ be a component, obtained by cutting along the maximal collection of normal $2$-spheres, which has one boundary component and which does not contain a vertex. Let $f:S^{2}\rightarrow M$ be an injective map whose image is the normal $2$-sphere $\delta M_{0}$. Since $M_{0}$ is a fake $3$-ball, the map $f$ is null homotopic. Let $F:S^{2}\times I\rightarrow M_{0}$ be a homotopy between $f$ and a constant map. For $t\in I$ let $f_{t}:S^{2}\rightarrow M_{0},f_{t}(s)=F(s,t)$. We can assume that for all but finitely many values $t,f_{t}$ meets the $1$-skeleton $T^{1}$ of the triangulation transversely and for each $t$ for which the map $f_{t}$ does not meet $T^{1}$ transversely, there is precisely one point where $f_{t}$ is tangential to $T^{1}$. Let $t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}$ be the values of $t$ for which $f_{t}$ does not meet $T^{1}$ transversely and put $t_{0}^{\prime}=0,t_{n+1}^{\prime}=1$. For $i=0,1,...,n$ choose $t_{i}\in(t_{i}^{\prime},t_{i+1}^{\prime})$ and put $f_{i}=f_{t_{i}}$.
@@
Let $W_{i}$ be the set of intersections of $f_{i}(S^{2})$ with the $1$-skeleton of $T$ and let $w_{i}=|W_{i}|$. Rearrange the weights $w_{i}$ into a finite non-increasing sequence. Order these sequences lexicographically. The width of $T$ is the minimum sequence of weights as $F$ ranges over all possible homotopies. If F realizes the width of $T$, then $F$ is said to be in thin position. Now revert to the original ordering of the $w_{i}$’s. Clearly for each $i$ either $w_{i+1}=w_{i}+2$ or $w_{i+1}=w_{i}-2$. Note that $w_{0}$ is the number of intersections of $\delta M_{0}$ with the $1$-skeleton. For $F$ in thin position, we are particularly interested in the values $i$ for which $w_{i-1}=w_{i+1}=w_{i}-2$. For such an $i,f_{i}(S^{2})$ is called a thick sphere, while if $w_{i-1}=w_{i+1}=w_{i}+2$, then $f_{i}(S^{2})$ is called a thin sphere. Our interest will be in the first thick sphere. Let $S=f_{k}(S^{2})$ be the first thick sphere.. Each of the preceding $f_{j}(S^{2})=S_{j}$, $1\leq j\leq k$, satisfies $w_{j-1}=w_{j}-2$, so that $W_{j}=W_{j-1}\cup\{u_{j},v_{j}\}$ where $u_{j},v_{j}$ are a pair of points, labelled $j,j$, from a particular $1$-simplex $\gamma_{j}$.
(a)(b)
Figure 7.
@@
Now consider $f_{k}:S^{2}\rightarrow M_{0}$ with $S=f_{k}(S^{2})$. We follow the argument of [5]. We know that the homotopy going from $f_{k}$ to $f_{k+1}$ results in the deletion of two points $x,y$ in a $1$-simplex $\gamma$ and the homotopy going from $f_{k}$ to $f_{k-1}$ deletes points $u,v$. We also know that these pairs must be + - pairs. All four points must belong to the same $2$-piece, for if there is no $2$-piece that contains all four points, then the order of the homotopies can be changed and the weight of the total homotopy reduced - one peak is replaced by two smaller peaks. Also if $x,y,u,v$ are in the same $2$-piece and $x,y$ are still in the same $2$-piece for the map $f_{k-1}$, then we can also swap the homotopies round and reduce weight. Similarly if $u,v$ are in the same $2$-piece for the map $f_{k+1}$, then we can swap the homotopies. It follows that the exceptional $2$-piece has two pairs of vertices that can be removed by a homotopy and that removing one pair disconnects the $2$-piece. The simplest such $2$-piece has boundary an $8$-track as in Figure 2 or Figure 7(a). Note that $x,y$ will be two points on one edge and ${u,v}$ will be the vertices on the opposite edge. Removing $u,v$ (labelled $k,k$) creates two 3-gons, with $x,y$ separated as shown in Figure 8. In an isotopy this is the only possibility for the exceptional $2$-piece. For a homotopy there are other possibilities such as in Figure 7(b).
@@
We will show that all the $2$-pieces apart from the exceptional one have no + - pairs of vertices which map to the same $1$-simplex. This is similar to the proof in [5]. Suppose $p,q$ are such a pair of vertices in a $2$-piece. These can be removed by a homotopy. Suppose first that $p,q$ are both different from any of $u,v,x,y$. Starting from the $(k-1)$-th stage, we can carry out a homotopy that removes $p,q$. We then add in the pair $u,v$ and remove $x,y$ and finally replace $p,q$ so that we are at the $(k+1)$-th stage. This sequence of homotopies has a lower sequence of weights than the original as the weight of the highest peak has been reduced by two. If, say $p=u$ but $q\not=v$, then we can remove $p,q$. This will not separate $x,y$ and so we can remove $x,y$ and then add in $u,v$. This has changed the points of $W_{k+1}$ by replacing $q$ by $v$ and this can be achieved by a homotopy which only affects the points on one $1$-simplex. Again this sequence of homotopies has a lower weight sequence.
@@
We see that $W_{0}$ is the set of vertices of a normal $2$-sphere $S_{0}$ parallel to the normal $2$-sphere $\delta M_{0}$. The pair $u_{1},v_{1}$ will be the ends of a returning arc in at least one $2$-simplex. In another $2$-simplex $u_{1},v_{1}$ will be joined by lines in both $S_{1}$ and $S$ to the vertices of an edge in $S_{0}$.
@@
Figure 9 shows the situation for $u_{1},v_{1}$ which lie in the $1$-simplex $\gamma_{1}$. In at least one $2$-simplex containing $\gamma_{1}$ as a face, $u_{1},v_{1}$ will be joined to the vertices $w,z$ of an edge of $S_{0}$, shown dashed red. In $P$, $w,z$ will be joined to points $x,y$ by lines, shown blue in Figure 9. Let $\nu_{1}$ be the permutation of $W$, which restricts to $(v_{1},x)(u_{1},y)$ for $W\cap\gamma_{1}$ and is the identity map on all the other $1$-simplexes. Under the corresponding homotopy the two lines become uncrossed if they were crossed before. Note that the end points are adjacent and stay adjacent because the homotopy takes place in $M_{0}$ and so the new lines of the homotopy do not intersect the lines of $S_{0}$. The pattern for $S_{0}$ in each $3$-simplex is as in Figure 1 and so the homotopy $\nu_{1}$ must take place in one of the regions which is not bounded by two parallel lines. This is illustrated in Figure 10, which indicates what can happen in the faces of a $3$-simplex. If a $2$-simplex $\sigma$ contains $\gamma_{1}$ as one of its faces, but has no intersection with $S_{0}$, then the points will be joined by a returning arc in $\sigma$.
@@
We have defined $\nu_{1}$. Let $\mu_{1}=\nu_{1}$. We now define $\nu_{i}$ and $\mu_{i}$ for $i=2,\dots,k$ to be permutations of $W$, where $\nu_{i}$ restricts to a permutation on the points of $W\cap\gamma_{i}$ and is the identity on the other points of $W$, and so that the homotopy associated with $\mu_{i}=\nu_{i}\mu_{i-1}$ moves the points of $W_{i}$ to the positions they should have in $P$. In at least one of the $2$-simplexes containing $\gamma_{i}$ as a face there are lines in $S$ joining $u_{j}$ to $u_{i}$ for some $j<i$ or there is a situation as in Figure 9, in which $u_{i},v_{i}$ are joined to the vertices of an edge of $S_{0}$.
If there are lines in $\mu_{i-1}S$ joining $u_{i}$ to $u_{j}$ and $v_{i}$ to $v_{j}$ it will be as in Figure 11(a) or there will be three labelled pairs joined by lines as in Figure 11(b). This is because $S_{i}$ has a returning arc joining $u_{i}$ and $v_{i}$ and this returning arc was created by one or more homotopy $f_{j}$ for $j<i$.
In Figure 11 (a) we see what happens in a $2$-simplex $\sigma_{i}$ containing $\gamma_{i}$, if the lines joining $u_{i}$ and $v_{i}$ come from a single pair $u_{j},v_{j}$. In Figure 11(b) the lines to $u_{i},v_{i}$ come from two pairs. The lines in blue are those of $\mu_{i-1}S\cap\sigma_{i}$. The lines in red are the ones we move them to in $\mu_{i}S$. The bottom edge is the $1$-simplex $\gamma_{i}$ containing the pair $u_{i},v_{i}$ which have labels $i,i$ in the labelling for $S$. In at least one of the $2$-simplexes containing $\gamma$ there will be a $1$-face different from $\gamma$ containing a pair of adjacent vertices with labels $l,l$ for $l<i$, which are joined by non-intersecting lines in $P$ to the pair $x,y$ in $\gamma_{i}$, or the situation will be as in Figure 11(b) so that there are pairs with labels less than $i$ in each of the other two faces. The induction step is to apply the permutation $\nu_{i}=(v_{i},x)(u_{i},y)$ to the set $\gamma_{i}\cap W$. Of course if $u_{i}=y$ we apply the transposition $(v_{i},x)$. Again with each transposition we also transpose all attaching lines in all $2$-simplexes containing $\gamma$. This change will not disturb any of the vertices of $W_{i-1}$ or the edges joining them. Note that $x,y$ will have labels bigger than $i$ as if, say, $x\in W_{i-1}$, then in $S_{i-1}$ there would be a point in the intersection of a $1$-simplex with $\mu_{i-1}S$ joined to both $u_{i}$ and $y$ and this does not happen in a spattern..
@@
Put $\mu=\mu_{k}$. We now know that $\mu S$ is a patterned surface. All the $2$-pieces, apart from the exceptional one, intersect each $1$-simplex at most once and so are $3$-sided or $4$-sided. The $8$-track shown in Figure 6(a) is the only possibility for the exceptional $2$-piece. Thus $S$ has become an almost normal $2$-sphere and we have a proof of the Poincaré Conjecture. In the homotopy $F$, the first thick sphere $S$ will also be the last thick sphere. After applying $\mu$, then for $j>k$ each step of the homotopy becomes an isotopy in which a returning arc joining adjacent points is removed.

### v2

@@ 2. Patterns in a Tetrahedron
A track is a connected pattern. If two patterns $P$ and $Q$ intersect each $1$-simplex in the same number of points then the patterns are said to be equivalent. Two equivalent disjoint tracks in the same $2$-complex are said to be parallel. We investigate tracks and patterns in a tetrahedron $T$, which we regard as the $2$-skeleton $|\rho^{2}|$ of a $3$-simplex $\rho$. We call a track in $T$ an $n$-track if it has $n$ intersections with edges.
@@
Figure 3 shows a $12$-track. A pattern in $T$ can only have two types of track. There can be $3$-tracks, each of which separates one of the corner vertices from the other three vertices. There can be one other parallel set of tracks each of which is a $4n$-track for the same positive integer $n$. Each such track separates the four vertices into two pairs. If $n>0$ is even, the $4n$-track separating $u,v$ from $w,z$ is as in Figure 2 but with $n-1$ lines crossing from $uv$ to $wz$. While if $n>1$ is odd the $4n$-track separating $u,z$ from $v,w$ is as in Figure 3 but with $n-1$ lines crossing from $uv$ to $wz$.
@@
Orient a track in $T$ by choosing a positive direction as one goes along the track. At adjacent points of intersection of a track with an edge the directions of the track will be opposite to each other. This gives what we call a + - pair, i.e. a pair of points - not necessarily adjacent - of intersection points of an edge with a track where the track has opposite directions. This will be of more significance for singular “tracks ” or stracks as shown in Figures 4 and 5, as if a pair of points of intersection on an edge can be removed by a homotopy, then the pair is a + - pair. A track in $T$ has a + - pair if and only if it is not a $3$-track or a $4$-track, since adjacent points of intersection of a track with an edge will be a + - pair.
@@
Let $f:S^{2}\rightarrow M$ be a general position map (see Hempel [3], Chapter 1), in which $f$ is in general position with respect to a triangulation $K$ of $M$. An $i$-piece of $f$ is defined to be a component of $f^{-1}(\sigma)$ where $\sigma$ is an $(i+1)$-simplex of $K$. Thus a 00-piece is a point of $S^{2}$. A $1$-piece is either an scc (simple closed curve) or an arc joining two 00-pieces. If there are no $1$-pieces that are scc’s, then each $2$-piece has boundary that is is a union of $1$-pieces. One can use surgery along simple closed curves to change $f$ to a map in which there are no $1$-pieces that are scc’s, and in which every $2$-piece is a disc. The $2$-pieces will then give a cell decomposition (tessellation) of the $2$-sphere.
@@
If $R$ is a $1$-piece with end points $u,v$ whose images under $f$ are in the same $1$-simplex, then the restriction of $f$ to $R$ is called a returning arc.
If $f:S^{2}\rightarrow M$ is a general position map with no $1$-pieces that are returning arcs, then the intersection of $f(S^{2})$ with the $2$-skeleton of $M$ is a spattern $sP$ and there is a homotopy from $f$ to $f^{\prime}:S^{2}\rightarrow M$ in which the image is the spatterned surface determined by $sP$.
Let $\gamma$ be a $1$-simplex of $M$. Two points $p,q\in\gamma\cap f(S^{2})$ are said to be removable if there is a homotopy from $f$ to a map $f^{\prime}:S^{2}\rightarrow M$ such that $f(x)=f^{\prime}(x)$ for every $x$ that is not in the interior of a simplex with $\gamma$ as a face and $\gamma\cap f^{\prime}(S^{2})$ is the same as $\gamma\cap f(S^{2})$ but with $p,q$ removed.
The pair of end points of a returning arc $R$ are removable by the following homotopy. Let $\sigma$ be the $2$-simplex of $K$ such that $f(R)\subset\sigma$. Let $V$ be a regular neighbourhood of $R$ in $S^{2}$. Let $V^{\circ}$ be the interior of $V$ regarded as a subspace of $V$. Let $\beta V$ be the boundary of $V$ regarded as a subspace of $S^{2}$, so that $\beta V=V-V^{\circ}$. Let $\gamma$ be the $1$-simplex containing the end points of $R$. The regular neighbourhood $V$ is a disc and $\beta V=\delta V$ is a simple closed curve in $S^{2}$. The union of all the $3$-simplexes that contain $\gamma$ is a closed ball and $f(\beta V)\subset B^{\circ}-\sigma$, which is contractible. Define $f^{\prime}:S^{2}\rightarrow M$ so that $f^{\prime}$ is continuous, $f^{\prime}$ and $f$ are the same when restricted to $S^{2}-V^{\circ}$, and $f^{\prime}(V)\subset B^{\circ}-\sigma$. Note that removing $p$ may create more $1$-pieces that are returning arcs or sccs, but the size of the intersection with the $1$-skeleton goes down by two. In the case in which we are interested intersections, in which sccs can be removed as above, the maps $f$ and $f^{\prime}$ are homotopic. This is illustrated in Figure 4, where it is shown how the removal of the ends of a returning arc will create returning arcs in two other simplexes that have the same $1$-face.
@@
There are also removable pairs of points if a strack intersects a $1$-simplex in a + - pair. In [1] there is a mistake on page 253 in the section on simplifying surface maps. It is incorrectly asserted there that any pair of points in the intersection of the boundary of a $2$-piece with a $1$-simplex can be removed by a homotopy. Suppose $Q$ is a $2$-piece and that $\gamma$ is a $1$-simplex for which $\delta Q\cap\gamma$ contains at least one + - pair. Let $s:S^{1}\rightarrow\delta Q$ be as in the definition of a strack. There will be at least one + - pair in $\gamma$ for which there is an arc $I=[p,q]\subset S^{1}$ such that $s(p)=u$ and $s(q)=v$ and $s(I)$ intersects $\gamma$ only in its end points $u,v$. Such a pair is removable. Thus if $\gamma$ contains a + - pair, then it contains a removable pair. Let $p,q$ be a removable pair as above. There will be a map $s^{\prime}:[p,q]\rightarrow Q$ which is close to $s$, for which $s^{\prime}(p)=u,s^{\prime}(q)=v$ and the image of the open interval $(p,q)$ is contained in the interior of the $3$-simplex $\rho$ containing $Q$. Any two maps of $I$ into the interior of $\rho$ are homotopic. In particular there will be such a map that is close to the interval $J$ in $\gamma$ joining $p,q$. We can adjust $f:S^{2}\rightarrow M$ by a homotopy so that the piece $Q$ contains the image of this map. Thus $f^{\prime}:S^{2}\rightarrow M$ is the same outside $Q$ and on the boundary of $Q$, but in the interior of $Q$ takes the arc close to $s$ to the arc $J$. Now $f$ and $f^{\prime}$ are homotopic. Another homotopy in a neighbourhood of $I$, will give a new map $f^{\prime\prime}$ in which the pair $p,q$ has been removed. This homotopy is similar to the one for removing a returning arc.
@@
In both cases, a removable pair can be removed without disturbing any other points of intersection with the $1$-skeleton.
@@
Figure 5 shows an $11$-strack in $T$ that has no + - pairs of points.
@@
Figure 6 shows a spattern in $T$ that is a union of a red 12-track and a blue 3-track, and its underlying pattern, which is also a $3$-track and a $12$-track. The underlying pattern for the strack of Figure 5 has two $4$-tracks. and one $3$-track.
@@
Given two finite subsets $F_{1},F_{2}$ of the closed interval $I=[0,1]$ and a bijection $\beta_{I}:F_{1}\rightarrow F_{2}$, there is a continuous map $\phi_{I}:I\rightarrow I$ which restricts to $\beta$ on $F_{1}$ and to the identity on $\{0,1\}$. There is a homotopy between $\phi_{I}$ and the identity map.
@@ 3. The Proof
Let $M$ be a fake 3-ball. Let $M_{0}$ be a component, obtained by cutting along the maximal collection of normal $2$-spheres, which has one boundary component and which does not contain a vertex. By Van Kampen, $M_{0}$ is simply connected, and so it is either a ball or a fake ball. In either case there is a homotopy between the boundary and the constant map. Let $f:S^{2}\rightarrow M$ be an injective map whose image is the normal $2$-sphere $\delta M_{0}$. Let $F:S^{2}\times I\rightarrow M_{0}$ be a homotopy between $f$ and a constant map. For $t\in I$ let $f_{t}:S^{2}\rightarrow M_{0},f_{t}(s)=F(s,t)$. We can assume that for all but finitely many values $t,f_{t}$ meets the $1$-skeleton $T^{1}$ of the triangulation transversely and for each $t$ for which the map $f_{t}$ does not meet $T^{1}$ transversely, there is precisely one point where $f_{t}$ is tangential to $T^{1}$. Let $t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}$ be the values of $t$ for which $f_{t}$ does not meet $T^{1}$ transversely and put $t_{0}=0,t_{n+1}=1$. For $i=1,...,n$ choose $t_{i}\in(t_{i}^{\prime},t_{i+1}^{\prime})$ and put $f_{i}=f_{t_{i}}$.
@@
Let $W_{i}$ be the set of intersections of $f_{i}(S^{2})$ with the $1$-skeleton of $T$ and let $w_{i}=|W_{i}|$. Rearrange the weights $w_{i}$ into a finite non-increasing sequence. Order these sequences lexicographically. The width of $T$ is the minimum sequence of weights as $F$ ranges over all possible homotopies. If F realises the width of $T$, then $F$ is said to be in thin position. Now revert to the original ordering of the $w_{i}$’s. Clearly for each $i$ either $w_{i+1}=w_{i}+2$ or $w_{i+1}=w_{i}-2$. Note that $w_{0}$ is the number of intersections of $\delta M_{0}$ with the $1$-skeleton. For $F$ in thin position, we are particularly interested in the values $i$ for which $w_{i-1}=w_{i+1}=w_{i}-2$. For such an $i,f_{i}(S^{2})$ is called a thick sphere, while if $w_{i-1}=w_{i+1}=w_{i}+2$, then $f_{i}(S^{2})$ is called a thin sphere. Our interest will be in the first thick sphere. Since there are no removable pairs in a normal surface, $w_{1}=w_{0}+2$. Let $S=f_{k}(S^{2})$ be the first thick sphere. Each of the preceding $f_{j}(S^{2})=S_{j}$, $1\leq j\leq k$, satisfies $w_{j-1}=w_{j}-2$, so that $W_{j}=W_{j-1}\cup\{u_{j},v_{j}\}$ where $u_{j},v_{j}$ are a pair of points, labelled $j,j$, from a particular $1$-simplex $\gamma_{j}$.
@@
Now consider $f_{k}:S^{2}\rightarrow M_{0}$ with $S=f_{k}(S^{2})$. We follow the argument of [5]. We know that the homotopy going from $f_{k}$ to $f_{k+1}$ results in the deletion a removable pair in a $1$-simplex $\gamma$ and the homotopy going from $f_{k}$ to $f_{k-1}$ results in the removal of another pair. Both pairs must belong to the same $2$-piece, for if there is no $2$-piece that contains both pairs, so that one pair is in one $2$-piece and the other removable pair is in another $2$-piece, then the order of the homotopies can be changed and the weight sequence of the total homotopy reduced - one peak is replaced by two smaller peaks. Also if both pairs are in the same $2$-piece and one pair is not separated by the removal of the other pair, then we can also swap the homotopies round and get a lower weight sequence. It follows that the exceptional $2$-piece has two removable pairs and that removing one pair disconnects the $2$-piece. The simplest such $2$-piece has boundary an $8$-track as in Figure 2. Removing one pair (labelled $k,k$) creates two 3-tracks, with the pair on the opposite edge separated as shown in Figure 7. In an isotopy this is the only possibility for the exceptional $2$-piece. For a homotopy there are other possibilities such as the $12$-strack in Figure 8. In this case removing the pair labelled $k,k$ gives a blue $3$-track and a red $7$-strack. In both cases, note that the lines joining the pair labelled $k,k$ connect to points on the other two edges of the $2$–simplex. They are neither returning arcs, nor do they connect to points on the same $2$-simplex. If they did, then removal of the pair labelled $k,k$ would not disconnect the strack.
@@
[…]
Figure 8.
We will show that all the $2$-pieces apart from the exceptional one have no removable pairs. This is similar to the proof in [5]. Suppose $p,q$ are a removable pair of 00-pieces in a $2$-piece different from the exceptional $2$-piece. Starting from the $(k-1)$-th stage, we can carry out a homotopy that removes $p,q$. We then carry out the homotopies $f_{k}$ and $f_{k+1}$ before replacing $p,q$. This sequence of homotopies has a lower sequence of weights than the original, as the weight of the highest peak has been reduced by two lower ones. If $p,q$ are in the exceptional $2$-piece but are also in $W_{k-1}$ or $W_{k+1}$, then we can also construct a sequence of homotopies with a lower weight sequence.
@@
We see that $W_{0}$ is the set of vertices of a normal $2$-sphere $S_{0}=\delta M_{0}$. The pair $u_{1},v_{1}$ will be the ends of a returning arc in at least one $2$-simplex. In another $2$-simplex $u_{1},v_{1}$ will be joined by lines in both $S_{1}$ and $S$ to the vertices of an edge in $S_{0}$. Having such a situation in a $2$-simplex is the only way that removing $u_{1},v_{1}$ will give a normal pattern.
@@
Figure 9 shows the situation for $u_{1},v_{1}$ which lie in the $1$-simplex $\gamma_{1}$. In at least one $2$-simplex containing $\gamma_{1}$ as a face, $u_{1},v_{1}$ will be joined to the vertices $w,z$ of an edge of $S_{0}$, shown dashed red. In $P$, $w,z$ will be joined to points $x,y$ by lines, shown blue in Figure 9. Let $\nu_{1}$ be the permutation of $W$, which restricts to $(v_{1},x)(u_{1},y)$ for $W\cap\gamma_{1}$ and is the identity map on all the other $1$-simplexes. Under the corresponding homotopy the two lines become uncrossed if they were crossed before. Note that the points $x,y$ are adjacent and stay adjacent in subsequent homotopies because the homotopy takes place in $M_{0}$ and so the new lines of the homotopy do not intersect the lines of $S_{0}$. The pattern for $S_{0}$ in each $3$-simplex is as in Figure 1 and so the homotopy $\nu_{1}$ must take place in one of the regions which is not bounded by two parallel lines. This is illustrated in Figure 10, which indicates what can happen in the faces of a $3$-simplex. If a $2$-simplex $\sigma$ contains $\gamma_{1}$ as one of its faces, but has no intersection with $S_{0}$, then the points will be joined by a returning arc in $\sigma$. In $\sigma$, a $4$-track has been replaced by two $3$-tracks. If instead of joining the ordered pair $u_{1},v_{1}$ to the vertices of the second edge, we joined $v_{1},u_{1}$ to those vertices we would get a $6$-strack. To get a subgraph of $P$ one must join up the points to give two $3$-tracks.
@@
We have defined $\nu_{1}$. Let $\mu_{1}=\nu_{1}$. We now define $\nu_{i}$ and $\mu_{i}$ for $i=2,\dots,k$ to be permutations of $W$, where $\nu_{i}$ restricts to a permutation on the points of $W\cap\gamma_{i}$ and is the identity on the other points of $W$, and so that the homotopy associated with $\mu_{i}=\nu_{i}\mu_{i-1}$ moves the points of $W_{i}$ to the positions they should have in $P$. In at least one of the $2$-simplexes containing $\gamma_{i}$ as a face there are lines in $S$ joining $u_{j}$ to $u_{i}$ for some $j<i$ or there is a situation as in Figure 9, in which $u_{i},v_{i}$ are joined to the vertices of an edge of $S_{j}$ for $j<i$. This is because there cannot be a returning arc joining $u_{i}$ and $v_{i}$ in every $2$-simplex containing $\gamma_{i}$, as if there were then $S_{i}$ would not be connected.
If there are lines in $\mu_{i-1}S$ joining $u_{i}$ to $u_{j}$ and $v_{i}$ to $v_{j}$ it will be as in Figure 11(a) or there will be three labelled pairs joined by lines as in Figure 11(b). This follows because $u_{i},v_{i}$ is a removable pair in $S_{i}$ and so it will have a returning arc joining $u_{i}$ and $v_{i}$ and this returning arc was created by one or more homotopy $f_{j}$ for $j<i$.
In Figure 11 (a) we see what happens in a $2$-simplex $\sigma_{i}$ containing $\gamma_{i}$, if the lines joining $u_{i}$ and $v_{i}$ come from a single pair $u_{j},v_{j}$. In Figure 11(b) the lines to $u_{i},v_{i}$ come from two pairs. The lines in blue are those of $\mu_{i-1}S\cap\sigma_{i}$. The lines in red are the ones we move them to in $\mu_{i}S$. The bottom edge is the $1$-simplex $\gamma_{i}$ containing the pair $u_{i},v_{i}$ which have labels $i,i$ in the labelling for $S$. In at least one of the $2$-simplexes containing $\gamma$ there will be a $1$-face different from $\gamma$ containing an adjacent pair with labels $l,l$ for $l<i$, which are joined by non-intersecting lines in $P$ to the pair $x,y$ in $\gamma_{i}$, or the situation will be as in Figure 11(b) so that there are pairs with labels less than $i$ in each of the other two faces. The induction step is to apply the permutation $\nu_{i}=(v_{i},x)(u_{i},y)$ to the set $\gamma_{i}\cap W$. Of course if $u_{i}=y$ we apply the transposition $(v_{i},x)$. Again with each transposition we also transpose all attaching lines in all $2$-simplexes containing $\gamma$. This change will not disturb any of the vertices of $W_{i-1}$ or the edges joining them. Note that $x,y$ will have labels bigger than $i$ as if, say, $x\in W_{i-1}$, then in $S_{i-1}$ there would be a point in the intersection of a $1$-simplex with $\mu_{i-1}S$ joined to both $u_{i}$ and $y$ and this does not happen in a spattern.. Put $\mu=\mu_{k}$.
For each labelled point of $W$ we can choose a line in $\mu S$ joining the point to a point of $W_{0}$ or a point with a smaller label. This will give us a connected subgraph of both $P$ and $\mu S$ that contains every point of $W$. This must mean that $P$ is connected. Since $S$ is a $2$-sphere and $\mu S$ is obtained from $S$ by permuting the points of $W$ and the lines joining them, $\mu S$ is also a $2$-sphere. But we now know $P$ and $\mu S$ have a common spanning subtree. Also $P$ and $\mu S$ have the same number of edges. This means that $P$ is also a $2$-sphere, as $P$ and $\mu S$ will have the same Euler characteristic. The addition of each edge not in a spanning tree results in a region being divided in two, and so there is one extra face for each such edge..
@@
We now know that $\mu S$ is a patterned surface. All the $2$-pieces, apart from the exceptional one, intersect each $1$-simplex at most once and so are $3$-sided or $4$-sided. The $8$-track shown in Figure 6(a) is the only possibility for the exceptional $2$-piece. Thus $S$ has become an almost normal $2$-sphere and we have a proof of the Poincaré Conjecture. In the homotopy $F$, the first thick sphere $S$ will also be the last thick sphere. After applying $\mu$, then for $j>k$ each step of the homotopy becomes an isotopy in which a returning arc joining adjacent points is removed. There is now a new labelling of $W$ in which every point receives a label $j$, where $n+1\leq j\leq n$. The pair of points labelled $j$ will be a removable pair in $f_{j}(S^{2})$.